A man bought 9 pounds of coffee and 6 pounds of butter for a total of $42.75
A day later, the prices had not changed and he bought 10 pounds of coffee and 14 pounds of butter for $64.00
Find the price per pound of the butter.''.

Answer by mananth(16949)

coffee x $/kg
butter y $/kg

9 x + 6 y = 42.75 .............1
10 x + 14 y = 64 .............2
Eliminate y
multiply (1)by -14
Multiply (2) by 6
-126 x -84 y = -598.5
60 x 84 y = 384
Add the two equations
-66 x = -214.5
/ -66
x = 3.25 OR 3 1/ 4
plug value of x in (1)
9 x + 6 y = 42.75
29.25 + 6 y = 42.75
6 y = 42.75 -29.25
6 y = 13.5
y = 2.25 OR 2 1/4
coffee 3 1/4 $/kg
butter 2 1/4 $/kg
m.ananth@hotmail.ca